the sum of three consecutive even numbers is 12. find the three numbers. (Hint: let the first number be x)
Answer by math-vortex(472)

Hi, there--
.
Use the hint. Let the first number be x.
The numbers are consecutive even numbers (like 6, 8, 10), so the second number is two more than the first, x+2.
The third number is two more than the second number, so it is x+2+2, or simply x+4.
.
We are told that the sum of the three numbers is 12. Writing this as an equation, we have
x + x+2 + x+4 = 12
.
Solve for x by combining like terms.
3x + 6 = 12
3x = 12 - 6
3x = 6
x = 2
.
The first number is 2, the second is 4, and the third is 6. When we add them together, we have 2+4+6=12. Perfect!
.
Hope this helps. Please feel free to email me if you have questions about this problem.
.
Ms.Figgy
math.in.the.vortex@gmail.com


the sum of three consecutive even numbers is 12. find the three numbers. (Hint: let the first number be x)